# NAAQA: A Neural Architecture for Acoustic Question Answering

### v2

@@ 3.3 Comparing CLEAR2 and DAQA
Acoustic scenes in DAQA are much longer on average than the ones in CLEAR2. This results in much bigger input spectrograms, and, in turns, much higher computational requirements and longer training time. Besides computational costs, the reason for producing shorter acoustic scenes in CLEAR2 was to produce a task that could be solved by a human. It would be unreasonable to ask, e.g., if the “flute plays before the sixth violin” in a 2 minute audio clip.
Finally, the original DAQA [undefaa] and, consequently, our reconstruction (DAQA) suffer from the same problem as the original CLEAR. The same elementary sounds are used in the training and test scenes. Although scenes are still different between training and test set, this may cause the models to “remember” the elementary sounds rather than extracting their properties. In CLEAR2, this problem was mitigated by using different elementary sounds for the training and test set.
@@ 4 Method
We first describe the original Visual FiLM architecture [undefz] that we use as baseline model, then the proposed modifications that lead to our NAAQA architecture and, finally, NAAQA with a MALiMo module.
@@ 4.1 Baseline model: Visual FiLM
Both the proposed NAAQA and Visual FiLM [undefz] share an overall common architecture which is depicted in Figure 2. Visual FiLM, that we will use as baseline model, is inspired by Conditional Batch Normalization architectures [undefaap] and achieved state of the art results on the CLEVR VQA task [undeff]. The network takes a visual scene and a text-based question as inputs and predicts an answer to the question for the given scene. The text-processing module uses $G$ unidirectional gated recurrent units (GRUs) to extract context from the text input (yellow area in Figure 2). The visual scene is processed by the convolutional module (blue area in the figure). The first step of this module is feature extraction (orange box), performed by a Resnet101 model [undefaae] pre-trained on ImageNet [undefaaq]. The extracted features are processed by a convolutional layer with batch normalization [undefaap] and ReLU [undefaar] activation followed by $J$ Resblocks illustrated in details in the red area in the figure. Unless otherwise specified, batch normalization and ReLU activation functions are applied to all convolutional layers. Each Resblock $j$ comprises convolutional layers with $M$ filters that are linearly modulated by FiLM layers through the two $M\times 1$ vectors $\boldsymbol{\beta}_{j}$ (additive) and $\boldsymbol{\gamma}_{j}$ (multiplicative). This modulation emphasizes the most important feature maps and inhibits the irrelevant maps given the context of the question. $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ are learned via fully connected layers using the text embeddings extracted by the text processing module as inputs (purple area in the figure). The affine transformation in the batch normalization before the FiLM layer is deactivated. The FiLM layer applies its own affine transformation using the learned $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ to modulate features. Several Resblocks can be stacked to increase the depth of the model, as illustrated in Figure 2. Finally, the classifier module is composed of a $1\times 1$ convolutional layer [undefaas] with $C$ filters followed by max pooling and a fully connected layer with $H$ hidden units and an output size $O$ equal to the number of possible answers (Gray in Figure 2). A softmax layer predicts the probabilities of the answers. In order to use the Visual FiLM as a baseline for our experiments, we extract a 2D spectro-temporal representation of the acoustic scenes as depicted at the bottom of Figure 2. The Resnet101 pre-trained extractor expects a 3 channels visual input but the spectro-temporal representation comprises only 1 channel. To work around this constraint, the spectro-temporal information is simply repeated 3 times thus creating a 3 channels input (only when using Resnet101 as feature extractor). This modified spectro-temporal representation is then fed to the model as if it was an image which is the simplest way to adapt the unmodified visual architecture to acoustic data. We call this architecture Visual FiLM Resnet101.
@@
Figure 2: Common Architecture. Two inputs: a spectro-temporal representation of an acoustic scene and a textual question. The spectro-temporal representation goes through a feature extractor (Parallel and Interleaved feature extractor detailed in Section 4.2.1 for NAAQA and Resnet101 pretrained on ImageNet for Visual FiLM) and then a serie of $J$ Resblocks that are linearly modulated by $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ (learned from the question input) via FiLM layers. Coordinate maps are inserted before convolution blocks with a pink border. The output is a probability distribution of the possible answers.
@@ 4.2 The proposed NAAQA architecture
In order to define the NAAQA architecture, we performed a number of modifications to the baseline architecture that will be described in the following sections.
@@ 4.2.1 Feature Extraction
As in Visual FiLM, the first step in the NAAQA model is feature extraction (orange box in Figure 2). The most obvious adaptation of Visual FiLM to acoustic data is to retrain the feature extraction module on the scenes from CLEAR2. In order to do this, we used three 2D convolutional layers, with $3\times 3$ kernels, stride $2\times 2$ and $N_{1}$, $N_{2}$, and $N_{3}$ filters respectively followed by a $1\times 1$ convolutions with $N_{4}$ filters. We will refer to this model as NAAQA 2D Conv.
@@
However, because acoustic signals present unique properties, we introduce two feature extraction modules that are specifically tailored to modeling sounds: the Parallel feature extractor (Figure 3a) processes time and frequency features independently in parallel pipelines; the Interleaved feature extractor (Figure 3b) captures time and frequency features in a single convolutional pipeline. In both cases, the feature extractor is trained end-to-end with the rest of the network and uses a combination of 1D convolutional filters to process a 2D spectro-temporal representation. The 1D filters process the time and frequency axis independently as opposed to the 2D filters typically used in image processing.
@@
The design of the Parallel feature extractor (Figure 3(a)) is inspired by the work of [undefaai] [undefaai] where 1D filters are used to capture time and frequency features separately. While [undefaai] time-frequency model includes only 1 time and 1 frequency convolution which are then concatenated together, our extractor stacks multiple 1D convolutions in two parallel pipelines. The time and frequency features are only fused at the end of both pipelines. This yields more complex features. The frequency pipeline (green in the figure) comprises a serie of $K$ frequency blocks. Each block is composed of a 1D convolution with $N_{K}$ $3\times 1$ kernels and $2\times 1$ strides followed by a $1\times 2$ maxpooling. With a stride larger than $1\times 1$, the convolution operation downsamples the frequency axis and the pooling operation downsamples the time axis. This downsampling strategy allows features in both parallel pipelines to be of the same dimensions. The time pipeline (yellow in the figure) is the same as the frequency pipeline except that the convolutional kernel operates along the time dimension and the pooling along the frequency dimension. The convolution kernel is $1\times 3$ and the pooling kernel is $2\times 1$. The activation maps of both pipelines are concatenated channel-wise and a representation combining both the time and frequency features is created using a $1\times 1$ convolution [undefaas] with $P$ filters and a stride of one. The feature maps dimensionality is either compressed or expanded depending on the number of filters $P$ in the $1\times 1$ convolution. We call the corresponding model NAAQA Parallel. The $1\times 1$ convolution can also be removed thus leaving it up to the next $3\times 3$ convolution to fuse the time and frequency features.
@@
The Interleaved feature extractor (Figure 3(b)) processes the input spectrogram in a single pipeline composed of a serie of $K$ interleaved blocks (purple in the figure). Each block comprises a $1\times 3$ time convolution with $N_{K}$ filters and stride $1\times 2$ followed by a $3\times 1$ frequency convolution with $N_{K}$ filters and stride $2\times 1$. A $1\times 1$ convolution with $P$ filters processes the output of the last block to either compress or expand its dimensionality. We call the corresponding model NAAQA Interleaved Time.
@@
For all extractors, the convolutions in the first block comprise $N_{1}$ convolutional filters and the number of filters is doubled after each block ($N_{i}$ = $2N_{i-1}$). More blocks (higher $K$) results in more down sampling of the feature maps shape which brings down the computational cost of the model.
@@ 4.2.2 Coordinate maps for acoustic inputs
The Visual FiLM model concatenates coordinate maps (CoordConv [undefaaj]) to the input of convolutional layers (pink border boxes in Figure 2) when tackling the VQA task. In the visual domain both axis of an image encode spatial information. Coordinate maps have, therefore, the same meaning in the $x$ or $y$-axis.
@@ 4.2.4 NAAQA with a MALiMo module
In MALiMo [undefaa], [undefaa] add a second set of FiLM layers that acts has an auxiliary controller. The controller uses the extracted acoustic features to further modulate the intermediate Resblocks. To evaluate the impact of MALiMo on CLEAR2, a MALiMo module was added to NAAQA. We refer to this configuration by appending MALiMo ctrl to the names introduced above. In our implementation of the module we replaced LSTMs with GRUs and adapted the inputs to the acoustic features that we study.
@@ 5 Experiments
We perform experiments to compare the effect on performance of our modification to the baseline model. Most of our experiments are performed on the proposed CLEAR2 data. We first investigate the performance of different feature extraction methods and compare them to the Visual FiLM Resnet101 feature extractor. Then, we show the effect of time and frequency coordinate maps at different levels of the model. Moreover, we perform an hyper-parameters ablation study to reduce the complexity of the model. We finally test the addition of a MALiMo module to our model. In order to demonstrate the generality of our results, we compare the performance of our model on the CLEAR2 and DAQA datasets.
@@ 5.1 Acoustic Pre-processing
The raw acoustic signal (sampled at 48 kHz for CLEAR2 and 16kHz for DAQA) is processed to create a 2D time-frequency representation using Mel scale [undefaat] spectrograms. After some experimentation we decided to use 64 Mel coefficients for both CLEAR2 and DAQA computed over samples weighted by a Hanning window. For CLEAR2 the window size was of 512 samples ($\sim$10.6 msec) whereas for DAQA we used windows of 400 samples ($\sim$25ms) as in [undefaa]. The time shift between consecutive windows (stride) was also optimized depending on the properties of the audio data. For CLEAR2 we found the best results for a time shift of 2048 samples ($\sim$42.7 msec). This is possible because of the properties of the sustained notes in the data set which vary slowly in time. Using such a long time shift allowed us to reduce more than ten folds the computational costs in our experiments. Because the DAQA dataset contains sounds that are more quickly varying, the same optimization is not possible. We tested with time shifts of 160 samples (10ms) and 1600 samples (100ms) observing a 5% drop in accuracy for the longer stride. All results based on CLEAR2 are reported with long window shift (long stride), with the exception of the comparison between short and long stride on both CLEAR2 and DAQA in Supplementary C.
Because the duration of the scenes in CLEAR2 is not constant, we pad the spectrogram along the time axis with zeros so that all spectrograms have the same dimensions ($1\times 64\times 418$) which corresponds to a maximum length of $\sim$17.9 sec. The power spectrum is normalized to the mean and standard deviation of the training data with the goal of speeding up convergence [undefaau].
@@ 5.2 Experimental conditions
Unless specified otherwise, the models presented in subsequent sections are trained on the CLEAR2 dataset which comprises 50 000 scenes and 4 questions per scene for a total of 200 000 records from which 140 000 (70%) are used for training, 30 000 (15%) for validation and 30 000 (15%) for test. The test set is generated using a different set of elementary sound which ensure that the network can not memorize the elementary sounds and therefore acts has a better benchmark for generalization. Optimisation technique and other training settings are further described in Supplementary A. The results are reported in terms of accuracy, that is the percentage of correct answers on the total. Since initialization of deep architectures has a profound impact on training convergence, we developed a python library torch-reproducible-block [^3] The link will be made available upon publication to control the model initial conditions and design reproducible experiments. To ensure the robustness of the results, each model is trained 5 times with 5 different random seeds.
@@ 5.3 Initial model configuration
TABLE III: Results on CLEAR2. The table shows number of parameters, average accuracy (%), and standard deviation over 5 repetitions of the training. Overall accuracy as well as question-kind dependent accuracy are reported. The different configurations are reported in the same order as they are presented in the paper. The most common answer is “Yes”.
@@ 6 Results and Discussion
Main results on the CLEAR2 data set are reported in Table III. It reports the complexity of the models in terms of number of parameters, the overall accuracy and accuracy dependent on question’s type. First we report two theoretical baselines given by random chance and majority class answers. Then we report results based on the baseline model Visual FiLM Resnet101 using the initial configuration described in Section 5.3. This architecture achieves the lowest accuracy of all tested models with 62.3%. This suggests that Mel spectrograms have a very different structure than visual scenes and, as expected, the pre-learned knowledge gathered in a visual context does not transfer directly to the acoustic context.
@@ 6.1.1 Feature Extraction
We then tested two versions of the Interleaved feature extractor (Figure 3(b)). Results show that the order of the 1D convolutions in each block has a significant impact on performance. When the first 1D convolution in each block is done along the frequency axis (NAAQA Interleaved Freq), the network reaches an overall accuracy of 67.2%. It performs especially poorly with questions related to count (42.5%), count instruments (46.5%) and notes (48.0%). The performance on position questions is also the lowest among all the extractors. When the order of the convolution is reversed (NAAQA Interleaved Time), useful information is better captured and the network reaches 78.0% of overall accuracy. A possible explanation relates to the nature of the sounds in the CLEAR2 dataset which mainly consists of sustained musical notes. The time dimension at short scales does not contain much information that helps identifying the individual sounds. At larger scales, however, the time axis contains information relative to the scene as a whole which is exploited by higher level layers (Resblocks) to model the connections between different sounds. Because its stride is greater than 1x1, each 1D convolution downsamples the axis on which it is applied. When the first convolution is a frequency convolution, the frequency axis of the resulting features is downsampled which reduces the useful information that can be captured by the time convolution that follows.

### v3

@@ 3.3 Comparing CLEAR2 and DAQA
Acoustic scenes in DAQA are much longer on average than the ones in CLEAR2. This results in much bigger input spectrograms, and, in turns, much higher computational requirements and longer training time.
Finally, the original DAQA [undefab] and, consequently, our reconstruction (DAQA) suffer from the same problem as the original CLEAR. The same elementary sounds are used in the training and test scenes. Although scenes are still different between training and test set, this may cause the models to “remember” the elementary sounds rather than extracting their properties. In CLEAR2, this problem was mitigated by using different elementary sounds for the training and test set.
@@ 4 Method
We first describe the original Visual FiLM architecture [undefaa] that we use as baseline model, then the proposed modifications that lead to our NAAQA architecture and, finally, NAAQA with a MALiMo module.
@@ 4.1 Baseline model: Visual FiLM
Both the proposed NAAQA and Visual FiLM [undefaa] share an overall common architecture which is depicted in Figure 2. Visual FiLM, that we will use as baseline model, is inspired by Conditional Batch Normalization architectures [undefaaq] and achieved state of the art results on the CLEVR VQA task [undeff]. The network takes a visual scene and a text-based question as inputs and predicts an answer to the question for the given scene. The text-processing module uses $G$ unidirectional gated recurrent units (GRUs) to extract context from the text input (yellow area in Figure 2). The visual scene is processed by the convolutional module (blue area in the figure). The first step of this module is feature extraction (orange box), performed by a Resnet101 model [undefaaf] pre-trained on ImageNet [undefaar]. The extracted features are processed by a convolutional layer with batch normalization [undefaaq] and ReLU [undefaas] activation followed by $J$ Resblocks illustrated in details in the red area in the figure. Unless otherwise specified, batch normalization and ReLU activation functions are applied to all convolutional layers. Each Resblock $j$ comprises convolutional layers with $M$ filters that are linearly modulated by FiLM layers through the two $M\times 1$ vectors $\boldsymbol{\beta}_{j}$ (additive) and $\boldsymbol{\gamma}_{j}$ (multiplicative). This modulation emphasizes the most important feature maps and inhibits the irrelevant maps given the context of the question. $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ are learned via fully connected layers using the text embeddings extracted by the text processing module as inputs (purple area in the figure). The affine transformation in the batch normalization before the FiLM layer is deactivated. The FiLM layer applies its own affine transformation using the learned $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ to modulate features. Several Resblocks can be stacked to increase the depth of the model, as illustrated in Figure 2. Finally, the classifier module is composed of a $1\times 1$ convolutional layer [undefaat] with $C$ filters followed by max pooling and a fully connected layer with $H$ hidden units and an output size $O$ equal to the number of possible answers (Gray in Figure 2). A softmax layer predicts the probabilities of the answers. In order to use the Visual FiLM as a baseline for our experiments, we extract a 2D spectro-temporal representation of the acoustic scenes as depicted at the bottom of Figure 2. The Resnet101 pre-trained extractor expects a 3 channels visual input but the spectro-temporal representation comprises only 1 channel. To work around this constraint, the spectro-temporal information is simply repeated 3 times thus creating a 3 channels input (only when using Resnet101 as feature extractor). This modified spectro-temporal representation is then fed to the model as if it was an image which is the simplest way to adapt the unmodified visual architecture to acoustic data. We call this architecture Visual FiLM Resnet101.
@@
Figure 2: Common Architecture. Two inputs: a spectro-temporal representation of an acoustic scene and a textual question. The spectro-temporal representation goes through a feature extractor (Parallel and Interleaved feature extractor detailed in Section 4.2.1 for NAAQA and Resnet101 pretrained on ImageNet for Visual FiLM) and then a serie of $J$ Resblocks that are linearly modulated by $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ (learned from the question input) via FiLM layers. Coordinate maps are inserted before convolution blocks that are illustrated with a pink border. The output is a probability distribution of the possible answers.
@@ 4.2 The proposed NAAQA architecture
To create the NAAQA architecture, we made modifications to the baseline architecture that will be described in the following sections. The code is available on GitHub [^5] https://github.com/NECOTIS/NAAQA-Acoustic-Question-Answering.
@@ 4.2.1 Feature Extraction
As in Visual FiLM, the first step in the NAAQA model is feature extraction (orange box in Figure 2). The most obvious adaptation of Visual FiLM to acoustic data is to retrain the feature extraction module on the scenes from CLEAR2. To do this, we used three 2D convolutional layers, with $3\times 3$ kernels, stride $2\times 2$ and $N_{1}$, $N_{2}$, and $N_{3}$ filters respectively followed by a $1\times 1$ convolutions with $N_{4}$ filters. We refer to this model as NAAQA 2D Conv.
@@
However, as acoustic signals present unique properties, we introduce two feature extraction modules that are specifically tailored to sounds: the Parallel feature extractor (Figure 3a) processes time and frequency features independently in parallel pipelines; the Interleaved feature extractor (Figure 3b) captures time and frequency features in a single convolutional pipeline. In both cases, the feature extractor is trained end-to-end with the rest of the network and uses a combination of 1D convolutional filters to process a 2D spectro-temporal representation. The 1D filters process the time and frequency axis independently as opposed to the 2D filters typically used in image processing.
@@
The design of the Parallel feature extractor (Figure 3(a)) is inspired by the work of [undefaaj] [undefaaj] where 1D filters are used to capture time and frequency features separately. While [undefaaj] time-frequency model includes only 1 time and 1 frequency convolution which are then concatenated together, our extractor stacks multiple 1D convolutions in two parallel pipelines. The time and frequency features are only fused at the end of both pipelines. This yields more complex features. The frequency pipeline (green in the figure) comprises a serie of $K$ frequency blocks. Each block is composed of a 1D convolution with $N_{K}$ $3\times 1$ kernels and $2\times 1$ strides followed by a $1\times 2$ maxpooling. With a stride larger than $1\times 1$, the convolution operation downsamples the frequency axis and the pooling operation downsamples the time axis. This downsampling strategy allows features in both parallel pipelines to be of the same dimensions. The time pipeline (yellow in the figure) is the same as the frequency pipeline except that the convolutional kernel operates along the time dimension and the pooling along the frequency dimension. The convolution kernel is $1\times 3$ and the pooling kernel $2\times 1$. The activation maps of both pipelines are concatenated channel-wise and a representation combining both the time and frequency features is created using a $1\times 1$ convolution [undefaat] with $P$ filters and a stride of one. The feature maps dimensionality is either compressed or expanded depending on the number of filters $P$ in the $1\times 1$ convolution. We name the corresponding model as NAAQA Parallel. The $1\times 1$ convolution can also be removed thus leaving it up to the next $3\times 3$ convolution to fuse the time and frequency features.
@@
The Interleaved feature extractor (Figure 3(b)) processes the input spectrogram in a single pipeline composed of a serie of $K$ interleaved blocks (purple in the figure). Each block comprises a $1\times 3$ time convolution with $N_{K}$ filters and stride $1\times 2$ followed by a $3\times 1$ frequency convolution with $N_{K}$ filters and stride $2\times 1$. A $1\times 1$ convolution with $P$ filters processes the output of the last block to either compress or expand its dimensionality. We name the corresponding model as NAAQA Interleaved Time.
@@
For all extractors, the convolutions in the first block comprise $N_{1}$ convolutional filters and the number of filters is doubled after each block ($N_{i}$ = $2N_{i-1}$). More blocks (higher $K$) gives a larger downsampling of the feature maps which brings down the computational cost of the model.
@@ 4.2.2 Coordinate maps for acoustic inputs
When tackling the VQA task, the Visual FiLM model concatenates coordinate maps (CoordConv [undefaak]) to the input of convolutional layers (pink border boxes in Figure 2). In the visual domain both axis of an image encode spatial information. Coordinate maps have, therefore, the same meaning in the $x$ or $y$-axis.
@@ 4.2.4 NAAQA with a MALiMo module
In MALiMo [undefab], [undefab] add a second set of FiLM layers that acts has an auxiliary controller. The controller uses the extracted acoustic features to further modulate the intermediate Resblocks. To evaluate the impact of MALiMo on CLEAR2, a MALiMo module was added to NAAQA. We refer to this configuration by appending MALiMo ctrl to the names introduced above. In our implementation of the module we replaced LSTMs with GRUs and adapted the inputs to the acoustic features that we study.
@@ 5 Experiments
We perform experiments to compare the effect on performance of our modification to the baseline model. Most experiments are conducted on the proposed CLEAR2 dataset. We first investigate different feature extraction methods and compare them to the Visual FiLM Resnet101 feature extractor. Then, we show the effect of time and frequency coordinate maps at different levels of the model. Moreover, we perform an hyper-parameters ablation study to reduce the complexity of the model. We finally test the addition of a MALiMo module to our model. To demonstrate the generality of the results, we compare the performance of our model on CLEAR2 and DAQA datasets.
@@ 5.1 Acoustic Pre-processing
The raw acoustic signal (sampled at 48 kHz for CLEAR2 and 16kHz for DAQA) is processed to create a 2D time-frequency representation with Mel scale [undefaau] spectrograms. After preliminary tests it was decided to extract 64 Mel coefficients for both CLEAR2 and DAQA computed over samples weighted by a Hanning window. The window size was of 512 samples ($\sim$10.6 msec) for CLEAR2 whereas for DAQA it was of 400 samples ($\sim$25ms) as in [undefab]. The time shift between consecutive windows (stride) was also optimized depending on the characteristics of audio data. We found that the best results for CLEAR2 was a time shift of 2048 samples ($\sim$42.7 msec). This is feasible because of the sustained notes which vary slowly in time. Using such a long time shift allowed us to reduce more than ten folds the computational costs. As DAQA contains sounds that are shorter and less stable, the same optimization is not feasible. In fact, with a time shift of 1600 samples (100ms) a 5% drop in accuracy is observed in comparison with 160 samples (10ms) shifts. All results based on CLEAR2 are reported with long window shifts (long stride), with the exception of the comparison between short and long strides on both CLEAR2 and DAQA in Supplementary Materials.
As duration of scenes are not constant in CLEAR2, spectrograms are zero padded along the time axis so that they all have the same dimension ($1\times 64\times 418$) which corresponds to a maximum length of $\sim$17.9 sec. The power spectrum is normalized to the mean and standard deviation of the training data with the goal of speeding up convergence [undefaav].
@@ 5.2 Experimental conditions
Unless specified otherwise, the models presented in subsequent sections are trained on the CLEAR2 dataset which comprises 50 000 scenes and 4 questions per scene for a total of 200 000 records from which 140 000 (70%) are used for training, 30 000 (15%) for validation and 30 000 (15%) for test. The test set is generated using a different set of elementary sounds which ensures that the network could not memorize them and can therefore acts has a better generalization benchmark. The optimization techniques and other training settings are further described in Supplementary Materials. Results are reported in terms of accuracy, that is in percentage of correct answers over the total. Since initialization of deep architectures has a profound impact on training convergence, we developed a python library torch-reproducible-block [^6] https://github.com/NECOTIS/torch-reproducible-block to control the model initial conditions and design reproducible experiments. To ensure the robustness of the results, each model is trained 5 times with 5 different random seeds.
@@ 5.3 Initial model configuration
TABLE III: Results on CLEAR2. Table gives the number of parameters, average accuracy (%), and standard deviation over 5 repetitions of the training. Overall accuracy as well as question-kind dependent accuracy are reported. Different configurations are reported in the same order as they are discussed in the paper. The most common answer is “Yes”.
@@ 6 Results and Discussion
Main results on the CLEAR2 data set are presented in Table III. The complexity of the models in terms of number of parameters, the overall accuracy and accuracy dependent on question’s type are reported. Results from two theoretical baselines - random chance and majority class answers - are first given. Then we report results from the Visual FiLM Resnet101 baseline model with the initial configuration described in Section 5.3. This architecture achieves the lowest accuracy of 62.3% in comparison with all tested models. As expected, the pre-learned knowledge gathered in a visual context does not transfer directly to the acoustic context. Mel spectrograms have a very different structure than visual scenes features.
@@ 6.1.1 Feature Extraction
We then tested two versions of the Interleaved feature extractor (Figure 3(b)). The computation order of the 1D convolutions in each block has a significant impact on performance. When the first 1D convolution in each block is computed along the frequency axis (NAAQA Interleaved Freq), the network reaches an overall accuracy of 67.2%. It performs especially poorly with questions related to count (42.5%), count instruments (46.5%) and notes (48.0%). The performance on position questions is also the lowest among all extractors. When the computation order of the convolution is reversed (NAAQA Interleaved Time), information is better captured and the network reaches 78.0% of overall accuracy. A possible explanation relates to the nature of the sounds in the CLEAR2 dataset which mainly consists of sustained musical notes. The time dimension at short scales does not contain much information that helps identifying the individual sounds. At larger scales, however, the time axis contains information relative to the scene as a whole which is exploited by higher level layers (Resblocks) to take into account the connections between different sounds. Because its stride is greater than 1x1, each 1D convolution downsamples the axis on which it is applied. When the first is a frequency convolution, the frequency axis of the resulting features is downsampled which reduces the information that can be captured by the time convolution that follows.